OpenStudy (anonymous):

find the domain and range of each function f(x)=log3x D:_________ R:_________ V.A_________ (V.A stand for vertical asymptote)

OpenStudy (helder_edwin):

any type of logarithm works only for positive numbers so \[ \large 3x>0\Rightarrow x>0/3=0 \] so the domain is \((0,+\infty)\)

OpenStudy (anonymous):

For the domain, you will be able to determine its value by considering that your base is positive.

OpenStudy (helder_edwin):

the logarithm of a number can be positive or negative so the range is \(\mathbb{R}\)

OpenStudy (helder_edwin):

a logarithmic function has a vertical asymptote when its argument is zero. so \[ \large 3x=0\Rightarrow x=0/3=0 \] therefore VA: x=0